OpenStudy (anonymous):

Joe has a collection of nickels and quarters that is worth $7.40. If the number of nickels were doubled and the number of quarters were increased by 8, the value of the coins would be $10.80. How many quarters does he have?

OpenStudy (anonymous):

I've already answered this one today, so you should be able to check my profile for the answer

OpenStudy (anonymous):

I just asked the same question......lol

OpenStudy (anonymous):

But I didn't understand the answer unfortunately

OpenStudy (anonymous):

Let x be the number of nickles, and let y be the number of quarters. Since a nickel is worth 5 cents, the total value of the nickels will be .05 times x, or .05x. Likewise, the total value of the quarters will be .25y. We know the total value of the collection of nickels and quarters is 7.40 so we can write out an equation: \[.05x+.25y = 7.40\] Does that make sense so far?

OpenStudy (anonymous):

(not done with the problem, just want to make sure we are on the same page)

OpenStudy (anonymous):

yes, I got it

OpenStudy (anonymous):

Joe has 24 quarters.

OpenStudy (anonymous):

alright, now lets look at the next part of the problem. It says the number of nickels was doubled. So if there were x nickels at first, now there will be 2x nickels. Looking at the quarters, it was increased by 8, so if there were y quarters at first, now there are (y + 8) quarters.

OpenStudy (anonymous):

Thank you crazy, but I need to figure out how to set it up and then solve it so that if I have to do it again, I know how instead of coming in here to ask.

OpenStudy (anonymous):

I got that part, it was after that I got confused

OpenStudy (anonymous):

5p+25q=740 5p(2)+25(q+8)=1080 10p+25q+200=1080 10p+50q=1480 10p+25q=880 25q=600 q=24 p=28 Joe has 24 quarters..

OpenStudy (anonymous):

what I nt get is how everyone is going from 10x+25y=800 and getting the answer

OpenStudy (anonymous):

Joe has 24 quarters.

OpenStudy (bahrom7893):

crazy go to your own thread

OpenStudy (anonymous):

So now using the same idea we did the first time, if each nickel is worth 5 cents, then the value of the nickels will be .05(2x) = x (because .05*2 = 1) The value of the quarters will be .25(y+8) = .25y+2 So the equation we get is \[x+.25y+2 = 10.80 \Rightarrow x+.25y = 8.80\]

OpenStudy (anonymous):

man deja vu all over again. i could have sworn i just did this

OpenStudy (anonymous):

You did, but I got lost and didn't understand your answer

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Now looking at the two equations: \[.05x+.25y=7.40\]\[x+.25y=8.80\] we need to find the solution. Noticing that there is a .25y in both equations, I will subtract them to make them cancel out. This is called Elimination. So eq 2 minus eq 1 gives: \[(1-.05)x = 8.80-7.40 \Rightarrow .95x = 1.40\]

OpenStudy (anonymous):

At which part are you lost at?

OpenStudy (anonymous):

from where you started the second part of the equation

OpenStudy (anonymous):

The\[x+.25y+2 = 10.80\] part?

OpenStudy (anonymous):

yes I dont get how the answer is gotten fromhere

OpenStudy (anonymous):

I think for me, that with the answers being so split up into different sections that I can't follow along with what your doing. I have to keep scrolling up to try to figure out what you had and what you did..........

OpenStudy (anonymous):

heres what I got so far 5p+25q=740 5p(2)+25(q+8)=1080 10p+25q+200=1080 10p+50q=1480 10p+25q=880 =?

OpenStudy (anonymous):

lmao, hi melissa it's diana

OpenStudy (anonymous):

lol i figured we were in the same class

OpenStudy (anonymous):

So for the second equation, I have twice as many nickels as I did before. So instead of 'x' nickels, i will have 2 times x, or 2x nickels. But each nickel is worth 5 cents. So to figure out how much money I have, I would multiply that 2x by .05, that gives : \[.05*2x = (.05*2)x = (1)x = x\] Thats where the x came from. @lilangl The reason it seems like we are spliting up the information is because these type of problems are called Systems of Equations. We have two things we dont know, the number of quarters and the nickels, so we need two equations to solve this. So once we create the first equation, put it to the side, and create another one.

OpenStudy (anonymous):

this site is a godsend

OpenStudy (anonymous):

o0o0o i get it now

OpenStudy (anonymous):

i still dont get number 4 though

OpenStudy (anonymous):

I don't mean that kind of split up, I just meant that part of the answer is about 15 sections up, then the next part is about 10 sections up, then another part 2 sections up............ instead of the answer being all part of the same post so that I can see how it breaks down.

OpenStudy (anonymous):

oh oh oh. my apologies on that . That happens a lot >.<

OpenStudy (anonymous):

I think I got that one right, let me figure this out and I'll let you know what I got

OpenStudy (anonymous):

can i take another crack at this or should i just shut up?

OpenStudy (anonymous):

I welcome any and all help satellite

OpenStudy (anonymous):

you know your input is valued satellite >.>

OpenStudy (anonymous):

well maybe it would just be more confusing than helpful

OpenStudy (anonymous):

I'm sorry I'm such an idiot when it comes to this, but generally, once I understand, I can do them by myself

OpenStudy (anonymous):

I would appreciate the help....im supposed to be leaving for a party soon-ish lolol

OpenStudy (anonymous):

lol, have fun for me Joe, I'll be stuck here trying to figure out how to do my graph problems next

OpenStudy (anonymous):

Sure thing :P

OpenStudy (anonymous):

already crying just thinking about it

OpenStudy (anonymous):

Did you have any luck figuring this out Mel?

OpenStudy (anonymous):

honestly i thought i had it but I came close I am trying to use substitution on the equations that joemath was using plus im stuck on number 4 u?

OpenStudy (anonymous):

I totally lost it on this one. I got so far and I am lost, and the answers we got here were a little to far spread out to figure out what was going on.............. Did you get my reply on your number 4?